# Adapt Tool Schemas to the Models

### v1

@@ Abstract
Small language models (SLMs) offer significant computational advantages for tool-augmented AI systems, yet they struggle with tool-use tasks, particularly in selecting appropriate tools and identifying correct parameters. A common failure mode is schema misalignment: models hallucinate plausible but non-existent tool names that reflect naming conventions internalized during pretraining but absent from the provided tool schema. Rather than forcing models to adapt to arbitrary schemas, we propose adapting schemas to align with models’ pretrained knowledge. We introduce PA-Tool (Pretraining-Aligned Tool Schema Generation), a training-free method that leverages peakedness—a signal from contamination detection indicating pretraining familiarity—to automatically rename tool components. By generating multiple candidates and selecting those with highest output concentration across samples, PA-Tool identifies pretrain-aligned naming patterns. Experiments on MetaTool and RoTBench show improvements of up to 17% points, with schema misalignment errors reduced by 80%. PA-Tool enables small models to approach state-of-the-art performance while maintaining computational efficiency for adaptation to new tools without retraining. Our work demonstrates that schema-level interventions can unlock the tool-use potential of resource-efficient models by adapting schemas to models rather than models to schemas.  [^1] We will release our code upon publication of the paper.
@@ 1 Introduction
Tool-augmented language models have become essential components of modern AI systems, with tool-agent frameworks being widely adopted across various applications Schick et al. (2023). As these systems mature, there is growing interest in deploying small language models (SLMs) due to computational constraints and the need for edge deployment Belcak et al. (2025); Chen et al. (2025b). In these systems, SLMs must perform two critical operations: tool selection (identifying which function to call) and parameter identification (providing appropriate arguments). However, small models struggle significantly with these tasks, exhibiting severe performance degradation particularly for models below 10B parameters that are most viable for edge deployment Patil et al. (2025); Erdogan et al. (2024).
@@
Figure 1: Effect of schema alignment on tool invocation. Top: Models learn tool documentation during pretraining. Bottom-Left: When schemas misalign with these internalized patterns, models generate plausible but non-existent tools. Bottom-Right: Schema alignment with pretrained knowledge prevents such errors.
A common failure pattern is schema misalignment: models hallucinate plausible but nonexistent components within the tool schema—a hierarchical structure consisting of tools and parameters, each with associated descriptions. As illustrated in Figure 1 (Top), models encounter various tool documentation patterns during pretraining, internalizing common naming conventions and API structures. However, when deployed with a specific tool schema, misalignment can occur. For instance (Figure 1, Bottom-Left), when the actual tool name to invoke is get_user_id, the model may instead generate get_customer_id—a similar but non-existent tool within the tool schema—leading to execution failure. Critically, these misaligned names are not random; they reflect naming patterns the model frequently encountered during pretraining but that are absent from the current tool schema. This raises a fundamental question: rather than forcing models to adapt to arbitrary tool schemas, what if we adapt the schemas to align with the naming conventions the models have internalized deeply? (Figure 1, Bottom-Right)
@@
We hypothesize that by renaming tools to match patterns in the model’s pretraining data—transforming get_user_id to get_customer_id if the latter better aligns with learned representations—we can reduce errors and improve tool-calling accuracy. To operationalize this idea, we propose PA-Tool (Pretraining-Aligned Tool Schema Generation), which generates a mapping between original tool names and pretrain-aligned alternatives. To identify such pretrain-aligned patterns, we leverage insights from contamination detection methods Dong et al. (2024), which detect whether models encountered specific data during pretraining. A key finding from these studies is that contaminated data—patterns frequently seen during training—exhibits peakedness: models generate highly similar outputs across multiple sampling attempts, creating a concentrated distribution. This occurs because repeated exposure to specific patterns during pretraining strengthens internal representations, reducing output variance around familiar patterns Shi et al. (2024a); Li (2023). We adopt this peakedness as a signal for the pretraining familiarity of tool schemas.
PA-Tool renames each component in a schema in three stages: (1) instructing the target language model to generate multiple candidate names based on a description of the component Chen et al. (2021), (2) computing the peakedness of each candidate, i.e., the number of other candidates that are sufficiently similar based on character-level edit distance Levenshtein (1965), and (3) selecting the candidate with the highest peakedness as the new name, as it is considered to best align with the model’s pretrained representation.
@@
Experiments on MetaTool and RoTBench with SLMs demonstrate PA-Tool’s effectiveness across diverse scenarios. On MetaTool, PA-Tool achieves substantial improvements across all subtasks, with the most significant gains when models must determine when no suitable tool exists (up to 17.0%) and when reasoning about multiple tools simultaneously (up to 9.6%). These improvements demonstrate PA-Tool’s effectiveness in addressing schema misalignment, which becomes particularly critical in complex tool selection scenarios. On RoTBench, PA-Tool maintains robust performance in both single-turn (5-10%) and multi-turn settings (4-10%), demonstrating that alignment benefits persist across extended conversational contexts. Notably, PA-Tool enables SLMs to approach the performance of closed-source models, with configurations like Qwen2.5-7B-Instruct matching or exceeding Claude Sonnet 4.5 on tool selection tasks.
Our comprehensive analysis reveals that PA-Tool’s effectiveness stems from directly addressing the predominant failure mode in SLMs. Schema misalignment errors—where models generate plausible but non-existent tool names—decrease by 80.0% with PA-Tool, while other error types show more modest reductions (18.8-24.0%). This targeted impact demonstrates that aligning schemas with pretrained knowledge resolves a bottleneck that disproportionately affects small models. Furthermore, PA-Tool’s training-free nature makes it highly practical for real-world deployment. This method requires only a simple one-time schema mapping without any model modification, enabling a model to use diverse tool sets without costly and complex retraining, and without having to worry about catastrophic forgetting or overfitting. This schema-level intervention represents an alternative approach: rather than forcing models to conform to arbitrary documentation standards, we adapt the interface to match the models’ internalized knowledge, enabling efficient tool use while preserving the computational advantages of small models.
@@
  • We propose PA-Tool, a training-free schema optimization method that repurposes peakedness—a contamination detection signal—to identify pretrain-aligned tool names.
  • We demonstrate improvements of up to 17% points across SLMs on MetaTool and RoTBench, with benefits extending from tool selection to parameter identification in both single-turn and multi-turn settings.
@@ 2.1 Emerging Agentic Frameworks
Recent advances in LLMs have enabled sophisticated agentic frameworks that decompose complex tasks into specialized modules, each coordinated by dedicated agents Shinn et al. (2023); Madaan et al. (2023); Wu et al. (2025); Agashe et al. (2025); Sapkota et al. (2026). In these multi-agent systems, SLMs are increasingly replacing larger models within individual modules to reduce computational costs and latency while maintaining specialized functionalities Belcak et al. (2025); Cheng et al. (2024). However, these systems remain vulnerable to cascading failures when SLM agents malfunction in foundational tool interaction tasks such as tool selection and parameter identification Erdogan et al. (2024); Patil et al. (2025). A particularly challenging failure mode is schema misalignment: SLM errors often manifest as plausible yet incorrect outputs—such as generating tool names that seem reasonable but don’t exist in the actual schema—making them difficult to detect and correct through traditional validation mechanisms. Our work addresses this vulnerability by generating pretraining-aligned schemas that improve SLM reliability in these critical operations.
@@ 2.2 Improving Tool Utilization in LLMs
As LLMs have demonstrated the capability to interact with external tools Schick et al. (2023); Hsieh et al. (2023), research has focused on evaluating and improving their tool-use capabilities. Evaluation efforts have developed benchmarks ranging from fine-grained assessments of tool selection and parameter identification Huang et al. (2024); Li et al. (2023); Ye et al. (2024); Chen et al. (2024); Patil et al. (2025) to end-to-end multi-step evaluation Seo et al. (2025); Yao et al. (2025); Trivedi et al. (2024); Qin et al. (2024); Shim et al. (2025).
Approaches to improving tool-use capabilities follow two main directions. Training-based methods employ supervised fine-tuning Qin et al. (2024); Liu et al. (2025); Zhang et al. (2025) or reinforcement learning Qian et al. (2025); Shi et al. (2024b); Feng et al. (2025); Chen et al. (2025a), but they require substantial data or computational resources. Training-free methods refine tool documentation Yuan et al. (2025); Qu et al. (2025) or leverage interaction histories Wang et al. (2024); Cui et al. (2025); Zhao et al. (2024); Fu et al. (2024). However, existing training-free approaches primarily focus on improving descriptions rather than aligning the schema itself with model-preferred representations. Our work directly addresses this gap by generating schemas aligned with pretraining distributions, targeting the root cause of schema misalignments.
@@ 2.3 Data Contamination Detection
Data contamination, defined as overlap between pretraining data and evaluation benchmarks, can inflate performance through memorization. Early detection methods used n-gram overlap Brown et al. (2020), but this approach fails to detect semantic paraphrasing. Probability-based approaches such as Min-k% Prob Shi et al. (2024a) and perplexity analysis Li (2023) require token probabilities that are unavailable in closed-source models. LLM Decontaminator Yang et al. (2023) uses auxiliary models for semantic similarity but introduces additional dependencies. Most relevant to our work, CDD Dong et al. (2024) identifies memorized patterns by measuring peakedness in sampled candidates, making it applicable to any black-box model. We adapt this peakedness mechanism to identify pretrain-aligned schema patterns, transforming contamination detection into a constructive tool for schema optimization.
@@ 3 PA-Tool : Pretraining-Aligned Tool Schema Generation
We now formalize our approach to generating pretrain-aligned tool schemas. A tool schema is hierarchical, structured documentation that defines tools and parameters along with detailed descriptions for each component. Our objective is to transform these descriptions into component names that the model is familiar with and that align with patterns the model has frequently encountered during pretraining. Specifically, we construct a dictionary mapping that associates each original component with a pretrain-aligned name—one that the language model would naturally generate based on its pretraining experience.
The core principle is straightforward: by analyzing the peakedness of a model’s output distribution when generating names from component descriptions, we can identify which naming patterns the model has most frequently encountered during pretraining. These pretrain-aligned names are hypothesized to improve the model’s ability to accurately invoke tools with correct parameters, reducing schema-related errors such as incorrect tool selections and parameter mismatches.
@@ 3.1 Framework Overview
Let $\mathcal{M}$ denote a pre-trained language model. Given a natural language description $d$ of a schema component’s functionality (e.g., a tool or a parameter description), our objective is to generate an optimal name $s^{*}$ that represents the naming pattern most deeply internalized by $\mathcal{M}$. As illustrated in Figure 2, our approach operates in three stages: (1) we provide $\mathcal{M}$ with a description of a component and instruct it to generate its name multiple times (Stage 1), (2) we compute peakedness scores that measure how many similar candidates cluster around each candidate name (Stage 2), and (3) we select the candidate with the highest peakedness as the pretrain-aligned name (Stage 3). By applying this process to each component in the schema hierarchy, we construct a dictionary mapping from original names to pretrain-aligned names. The complete algorithm is provided in Appendix A.
@@ 3.2 Stage 1: Candidate Generation
In this stage, we collect diverse candidate names of a component that the model may have encountered during the pretraining. As illustrated in Figure 2-1, given a component’s description, we sample $N$ candidate names $\mathcal{C}=\{s_{1},s_{2},\ldots,s_{N}\}$ from the language model with temperature $t\in(0,1]$. This temperature-controlled sampling explores the model’s learned distribution beyond the single greedy path, revealing diverse naming patterns Chen et al. (2021). The detailed prompt structure is provided in the Appendix B. Additionally, we generate a reference name $s_{\text{ref}}$ using greedy decoding ($t=0$) for tie-breaking purposes in the selection stage.
@@ 3.4 Stage 3: Representative Name Selection
This approach rests on the hypothesis that frequently occurring patterns in training data create local maxima in the model’s output distribution. When a naming pattern appears multiple times during training, the model develops stronger internal representations, leading to higher probability mass and reduced variance around these patterns. By identifying regions of high peakedness, we effectively locate these memorized naming conventions, which represent the most natural and well-formed component names according to the model’s learned knowledge. Through iterative application across all schema components (Figure 2-3), we obtain the final pretrain-aligned schema. An example of a schema generated through this process is provided in Appendix C.
@@ 4 Experimental Setup
PA-Tool 50.0 58.8 86.2 72.6 6.7 1.9 7.1 4.3
Qwen2.5-7B Base 59.6 74.4 78.3 78.3 49.5 20.0 8.6 5.7
Greedy 60.6 75.4 85.4 82.7 50.5 21.0 14.3 8.6
PA-Tool 64.1 78.4 88.2 84.9 55.2 21.9 18.6 11.4
@@
Table 1: Performance comparison on MetaTool and RoTBench. All metrics are reported as accuracy (%). Tool Sel. and Param Iden. denote Tool Selection and Parameter Identification, respectively. Bold indicates the best performance within each closed-source model and among all open-source models.
To evaluate whether PA-Tool improves models’ core tool-use capabilities, we conduct comprehensive experiments assessing two fundamental aspects: (1) Tool selection, whether models correctly identify the appropriate tool for a given query from available options, and (2) Parameter identification, whether models accurately extract necessary arguments to construct valid tool calls.

### v2

@@ Abstract
Small language models (SLMs) enable scalable multi-agent tool systems where multiple SLMs handle subtasks orchestrated by a powerful coordinator. However, they struggle with tool-use tasks, particularly in selecting appropriate tools and identifying correct parameters. A common failure mode is schema misalignment: models hallucinate plausible but nonexistent tool names that reflect naming conventions internalized during pretraining but absent from the provided tool schema. Rather than forcing models to adapt to arbitrary schemas, we propose adapting schemas to align with models’ pretrained knowledge. We introduce PA-Tool (Pretraining-Aligned Tool Schema Generation), a training-free method that leverages peakedness, a signal from contamination detection indicating pretraining familiarity, to rename tool components. By generating multiple candidates and selecting those with the highest peakedness across samples, PA-Tool identifies pretraining-aligned naming patterns. Experiments on MetaTool and RoTBench show improvements of up to 17%, with schema misalignment errors reduced by 80%. PA-Tool enables small models to approach state-of-the-art performance while maintaining computational efficiency in adapting to new tools without retraining. Our work demonstrates that schema-level interventions can unlock the tool-use potential of resource-efficient models by adapting schemas to models rather than models to schemas.  [^1] We will release our code upon publication of the paper.
@@ 1 Introduction
Tool-augmented language models have become essential components of modern AI systems Qu et al. (2025b). As these systems mature, there is growing interest in deploying small language models (SLMs, typically <<10B), driven by both multi-agent systems where a powerful model orchestrates reasoning while multiple SLMs handle subtasks Belcak et al. (2025) and edge deployment scenarios with strict resource constraints Chen et al. (2025b). In these systems, SLMs must perform two critical operations: tool selection (identifying which API to call) and parameter identification (providing appropriate arguments). However, SLMs struggle significantly with these tasks, exhibiting severe performance degradation compared to larger models Patil et al. (2025); Erdogan et al. (2024).
@@
Figure 1: Effect of schema alignment on tool invocation. Top: Models learn tool schema during pretraining. Bottom-Left: When schemas misalign with these internalized patterns, models generate plausible but nonexistent tools. Bottom-Right: Schema alignment with pretrained knowledge prevents such errors.
A common failure pattern is schema misalignment: even when appropriate tools are provided in context, models fail to invoke them and instead hallucinate plausible but nonexistent components (Figure 1, Bottom-Left) within the tool schema—a hierarchical structure of tools and parameters with descriptions. This suggests models are confusing unfamiliar tool schemas with familiar naming conventions internalized during pretraining. Therefore, we hypothesize that aligning tool schemas with patterns from the model’s internalized knowledge can reduce errors and improve tool-calling accuracy (Figure 1, Bottom-Right).
@@
To operationalize this idea, we propose PA-Tool (Pretraining-Aligned Tool Schema Generation), which generates a mapping between original tool component names and model-familiar alternatives. To identify what naming conventions the model is familiar with, we leverage insights from contamination detection methods Dong et al. (2024), which detect whether models encountered specific data during pretraining. A key finding from these studies is that patterns frequently seen during training exhibit peakedness: models generate highly similar outputs across multiple sampling attempts, creating a concentrated distribution Shi et al. (2024a); Li (2023). We adopt this peakedness as a signal for the model’s familiarity with tool names. Concretely, PA-Tool renames each component through three stages: (1) instructing the target language model to generate multiple candidate names based on a description of the component Chen et al. (2021), (2) computing the peakedness of each candidate, measured as the number of other candidates that are sufficiently similar based on character-level edit distance Levenshtein (1965), and (3) selecting the candidate with the highest peakedness as the new name, as it is considered to best align with the model’s pretrained knowledge.
@@
Experiments on MetaTool and RoTBench demonstrate PA-Tool’s effectiveness across diverse scenarios. On MetaTool, PA-Tool achieves improvements up to 17% when models must identify that no suitable tool exists and 9.6% for multi-tool selection scenarios. On RoTBench, performance gains of 5-10% (single-turn) and 4-10% (multi-turn) show that alignment benefits persist across extended contexts. Notably, PA-Tool enables SLMs like Qwen2.5-7B-Instruct to match or exceed Claude Sonnet 4.5 on tool selection tasks.
Our comprehensive analysis reveals that PA-Tool’s effectiveness stems from directly addressing the predominant failure mode in SLMs. Schema misalignment errors, where models generate plausible but nonexistent tool names, decrease by 80.0% with PA-Tool, while other error types show more modest reductions (18.8-24.0%). This targeted impact demonstrates that aligning schemas with pretrained knowledge resolves a bottleneck that disproportionately affects small models. We further validate that peakedness reliably captures learning exposure by showing it consistently increases as models encounter tool schemas during training. PA-Tool also demonstrates broad compatibility with existing approaches: it provides complementary gains when combined with retrieval-based correction, constrained generation, and description enhancement methods, and remains effective even after supervised fine-tuning. Notably, PA-Tool alone outperforms fine-tuned models on several subtasks without any training, while fine-tuning risks catastrophic forgetting to specific tool distributions. Beyond tool selection benchmarks, PA-Tool improves end-to-end task completion on $\tau$-Bench, confirming that schema alignment translates to practical performance gains.
This training-free approach requires only a one-time schema mapping: rather than forcing models to conform to arbitrary documentation, we adapt the interface to match their internalized knowledge, enabling diverse tool use without model modification, retraining, or catastrophic forgetting.
@@
  • We propose PA-Tool, a training-free schema optimization method that identifies pretraining-aligned tool names by repurposing peakedness, a signal from contamination detection.
  • We demonstrate improvements of up to 17% across SLMs on MetaTool and RoTBench, with benefits extending from tool selection to parameter identification in both single-turn and multi-turn settings.
@@ 2.1 Emerging Agentic Frameworks
Recent advances in LLMs have enabled sophisticated agentic frameworks that decompose complex tasks into specialized modules, each coordinated by dedicated agents Shinn et al. (2023); Madaan et al. (2023); Wu et al. (2025); Agashe et al. (2025); Sapkota et al. (2026). In these multi-agent systems, SLMs are increasingly replacing larger models within individual modules to reduce computational costs and latency while maintaining specialized functionalities Belcak et al. (2025); Cheng et al. (2024). However, these systems remain vulnerable to cascading failures when SLM agents malfunction in foundational tool interaction tasks such as tool selection and parameter identification Erdogan et al. (2024); Patil et al. (2025). A particularly challenging failure mode is schema misalignment: SLM errors often manifest as plausible yet incorrect outputs, such as generating tool names that seem reasonable but don’t exist in the actual schema, making them difficult to detect and correct through traditional validation mechanisms. Our work addresses this vulnerability by generating pretraining-aligned schemas that improve SLM reliability in these critical operations.
@@ 2.2 Improving Tool Utilization in LLMs
As LLMs demonstrate the capability to interact with external tools Schick et al. (2023); Hsieh et al. (2023), research has focused on evaluating and improving their tool-use capabilities. Evaluation efforts have developed benchmarks ranging from fine-grained assessments of tool selection and parameter identification Huang et al. (2024); Li et al. (2023a); Ye et al. (2024); Chen et al. (2024); Patil et al. (2025) to end-to-end multi-step evaluation Seo et al. (2025); Yao et al. (2025); Trivedi et al. (2024); Qin et al. (2024); Shim et al. (2025).
Approaches to improving tool-use capabilities follow two main directions. Training-based methods employ supervised fine-tuning Qin et al. (2024); Liu et al. (2025); Zhang et al. (2025) or reinforcement learning Qian et al. (2025); Zhou et al. (2025); Shi et al. (2024b); Feng et al. (2025); Chen et al. (2025a), but they require substantial data or computational resources. Training-free methods refine tool documentation Yuan et al. (2025); Qu et al. (2025a) or leverage interaction histories Wang et al. (2024); Cui et al. (2025); Zhao et al. (2024); Fu et al. (2024). However, existing training-free approaches primarily focus on improving descriptions rather than aligning the schema itself with model-preferred representations. Our work directly addresses this gap by generating schemas aligned with pretraining distributions, targeting the root cause of schema misalignments.
@@ 2.3 Data Contamination Detection
Data contamination, defined as overlap between pretraining data and evaluation benchmarks, can inflate performance through memorization. Early detection methods use n-gram overlap Brown et al. (2020), but these approaches fail to detect semantic paraphrasing. Probability-based approaches such as Min-k% Prob Shi et al. (2024a) and perplexity analysis Li (2023) require token probabilities that are unavailable in closed-source models. LLM Decontaminator Yang et al. (2023) uses auxiliary models for semantic similarity but introduces additional dependencies. Most relevant to our work, CDD Dong et al. (2024) identifies memorized patterns by measuring peakedness in sampled candidates, making it applicable to any black-box model. We adapt this peakedness mechanism to identify pretraining-aligned schema patterns, transforming contamination detection into a constructive tool for schema optimization.
@@ 3 PA-Tool : Pretraining-Aligned Tool Schema Generation
We now formalize our approach to generating pretraining-aligned tool schemas. A tool schema is structured documentation that defines tools and parameters along with descriptions for each component. Our objective is to rename these components with names that align with patterns the model has frequently encountered during pretraining.
@@ 3.1 Framework Overview
Let $\mathcal{M}$ denote a pre-trained language model. Given a natural language description $d$ of a schema component’s functionality (e.g., a tool or a parameter description), our objective is to generate an optimal name $s^{*}$ that represents the naming pattern most deeply internalized by $\mathcal{M}$. As illustrated in Figure 2, our approach operates in three stages: (1) we provide $\mathcal{M}$ with a description of a component and instruct it to generate its name multiple times (Stage 1), (2) we compute peakedness scores that measure how many similar candidates cluster around each candidate name (Stage 2), and (3) we select the candidate with the highest peakedness as the pretraining-aligned name (Stage 3). By applying this process to each component in the schema hierarchy, we construct a dictionary mapping from original names to pretraining-aligned names. The complete algorithm is provided in Appendix A.
@@ 3.2 Stage 1: Candidate Generation
In this stage, we collect diverse candidate names of a component that the model may have encountered during pretraining. As illustrated in Figure 2-1, given a component’s description, we sample $N$ candidate names $\mathcal{C}=\{s_{1},s_{2},\ldots,s_{N}\}$ from the language model with temperature $t\in(0,1]$. This temperature-controlled sampling explores the model’s learned distribution beyond the single greedy path, revealing diverse naming patterns Chen et al. (2021). The detailed prompt structure is provided in Appendix B. Additionally, we generate a reference name $s_{\text{ref}}$ using greedy decoding ($t=0$) for tie-breaking purposes in the selection stage.
@@ 3.4 Stage 3: Representative Name Selection
This approach rests on the hypothesis that frequently occurring patterns in training data create local maxima in the model’s output distribution. By identifying regions of high peakedness, we effectively locate these memorized naming conventions, which represent the most natural and well-formed component names according to the model’s learned knowledge. Through iterative application across all schema components (Figure 2-3), we obtain the final pretraining-aligned schema. An example of a schema generated through this process is provided in Appendix C. We repeat our three-stage process for all tool sets within the system. When name collisions occur across different tools due to highly similar descriptions, we resolve them using the priority-based locking mechanism (Appendix I).
@@ 4 Experimental Setup
MostFreq 50.5 56.3 84.3 73.0 12.4 9.5 5.7 2.9
Human 54.6 62.9 80.6 72.6 - - - -
PA-Tool 50 58.8 86.2 72.6 6.7 1.9 7.1 4.3
Qwen2.5-7B Base 59.6 74.4 78.3 78.3 49.5 20 8.6 5.7
Greedy 60.6 75.4 85.4 82.7 50.5 21 14.3 8.6
MostFreq 65.3 75.5 90.0 86.1 50.5 23.8 12.9 10
Human 65.8 76.7 78.6 82.1 - - - -
PA-Tool 64.1 78.4 88.2 84.9 55.2 21.9 35.7 11.4
@@
Table 1: Performance comparison on MetaTool and RoTBench. All metrics are reported as accuracy (%). Bold indicates the best performance within each open-source model and among all closed-source models.
@@ 4.1 Benchmarks
MetaTool. MetaTool (Huang et al., 2024) evaluates tool selection capabilities across 4,287 test cases in four subtasks: (1) Similar tests semantic comprehension by distinguishing tools with overlapping functionalities (e.g., Sudoku vs. Tic-Tac-Toe); (2) Scenario selects appropriate tools based on user-specific contexts and requirements (e.g., software engineers, students); (3) Reliability assesses whether models can directly indicate when no suitable tool is available rather than hallucinating; and (4) Multi-tool measures whether models can correctly select multiple tools when tasks require composition of functionalities.
RoTBench. RoTBench (Ye et al., 2024) evaluates whether models can robustly use tools when descriptions contain various noise perturbations (e.g. reversed names). However, we focus exclusively on the Clean environment to isolate the effects of PA-Tool on core tool-use capabilities rather than noise robustness. We evaluate on 105 test cases across 568 tools in both single-turn and multi-turn settings. RoTBench measures two fundamental tool use capabilities: (1) Tool Selection evaluates whether models correctly identify the appropriate tool; and (2) Parameter Identification assesses whether models accurately extract the required parameter set, conditioned on correct tool selection.
4.2 Models and Baselines
We focus on models with 8B parameters or fewer (SLMs), evaluating four open-source models: Qwen 2.5 3B/7B (Qwen et al., 2025), Llama-3.1 8B (Meta AI, 2024a), and Llama 3.2 3B (Meta AI, 2024b). For each model, we evaluate five configurations: (1) Base uses the original schema without modifications; (2) Greedy uses greedy decoding to generate deterministic schema names (no schema-change control); (3) MostFreq selects the most frequently generated candidate from our sampling process; (4) Human uses schemas where two PhD-level researchers with software engineering backgrounds manually renamed tools to more intuitive names (evaluated only on MetaTool due to RoTBench’s large tool-sets); (5) PA-Tool applies our proposed peakedness-based method. To contextualize SLM performance and quantify PA-Tool’s improvements, we also compare against three state-of-the-art closed-source models: GPT-4.1-mini (OpenAI, 2024), Gemini 2.5 Flash (Basu Mallick et al., 2025), and Claude Sonnet 4.5 (Anthropic, 2025).
4.3 Implementation Details
When generating schemas with PA-Tool, we use 32 candidates at temperature 0.4 with $\alpha=0.2$ (sensitivity analysis in Appendix J). Each component’s description is provided individually (See Figure 2-1; prompt template in Appendix B).
For benchmark inference, we use temperature 0 to ensure reproducible results across all experiments. We use accuracy as the primary evaluation metric across both benchmarks, measuring the percentage of test cases where the model’s predictions exactly match the ground-truth labels. Detailed evaluation protocols are in Appendix D.
5 Main Results
Table 1 presents comprehensive results across MetaTool and RoTBench benchmarks, demonstrating PA-Tool’s effectiveness.